OpenStudy (anonymous):

The price of a shirt was increased from $20.00 to $30.00. What was the percent of increase? SHOW WORK

OpenStudy (anonymous):

(new-old)/old (30-20)/20 50% increase

jhonyy9 (jhonyy9):

100 ... 20 x .... (30-20)=10 -------------- x= (10)(100)/20 = 50 % increase - hence this is clearly ?